Making a metric ton of paper requires 1100 kWh of power whereas making a metric ton of polystyrene requires about 280 kWh. Given taht an average 8oz cup weighs 10.1 g and an average polystyrene cup weighs 1.5g, what is the ratio of the power requirements on a per-cup basis? (paper/plastic)

for the paper cups, calculate the # paper cups/ton, then energy/paper cups. for poly cups calculate the same quantities. Take ratio
Answer #1

1 metric tons = 1000 kg = 106 g

Thus, power requirement for 106 g of paper production = 1.1*106 W-hr.

Thus, power requirement for the production of 10.1 g of paper = 1.1*10.1 = 11.11 W-hr/cup

power requirements for 106 g of polystyrene production = 0.28*106 W-hr

power requirement for the production of 1.5 g of polystyrene = 0.28*1.5 = 0.42 W-hr/cup

Thus, ratio of power requirements on cup basis = 11.11/0.42 = 26.45:1